.As a venture capitalist, you are considering investing in a startup company called Xevo. Xevo, like most startups, is not expected to generate profits in the near future. After careful analysis, you and your colleagues determine that Xevo will start generating profits of $400,000 at the end of the 6 th year. Thereafter, it will grow at 11% per year forever. Using a discount rate of 18%, and assuming you will be acquiring 50% of the company, what is the amount your venture capital firm would be willing to invest? (a) $2,857,143 (b) $1,248,883 (c) $1,695,575 (d) $1,058,376

Xevo is hot in the Valley, and there are other venture capitalists bidding for the 50% share of the company. Another venture capitalist, a long lost friend with whom you once took Corporate Finance together at OSU, says that it’s more likely that Xevo will be able to generate $100,000 in profits at the end of each year for the first 5 years. She agrees with you on Xevo’s prospects after that as well as on the discount rate. What’s her price tag for the 50% share in Xevo? (a) $1,214,735 (b) $1,405,242 (c) $1,561,600 (d) $2,857,143
Answer #1

(1):Here profit in year 7 = profit in year 6*(1+11%)

= 400,000*1.11

= 444,000. This will henceforth grow at 11% per year till infinity. Thus the value of this at the end of 6th year = 444,000/(18%-11%) = 6,342,857.14

Total value = 400,000/1.18^6 + 6,342,857.14/1.18^6

= 2,497,766.95

As the venture capital is acquiring 50% it will invest 2,497,766.95/50% = 1,248,883

Hence the answer is option “b” $1,248,883

(2): In this case value of Xevo will be =100,000/1.18 + 100,000/1.18^2 + 100,000/1.18^3 + 100,000/1.18^4 + 100,000/1.18^5 + 400,000/1.18^6 + 6,342,857.14/1.18^6

= 312,717.10+400,000/1.18^6 + 6,342,857.14/1.18^6

= 312,717.10+2,497,766.95

= 2,810,484.05

Thus price tag for 50% share here = 2,810,484.05/50% = 1,405,242

Thus the answer is option “b” $1,405,242